OpenStudy (anonymous):

. In solving the equation (x + 4)(x – 7) = -18, Eric stated that the solution would be x + 4 = -18 => x = -22 or (x – 7) = -18 => x = -11 However, at least one of these solutions fails to work when substituted back into the original equation. Why is that?

hero (hero):

Eric solved the problem incorrectly

OpenStudy (anonymous):

oh dear you can't do that, you have to get 0 on one side of the equation

OpenStudy (anonymous):

hero do you want to take this one?

hero (hero):

He should multiply the binomials together first, then add 18 to both sides. Then solve for x

hero (hero):

Eric should do this: (x + 4)(x – 7) = -18 x^2 - 3x - 28 = -18 x^2 - 3x - 10 = 0 x^2 - 5x + 2x -10 = 0 x(x-5)+2(x-5) = 0 (x-5)(x-2) = 0 x =2,5

OpenStudy (anonymous):

yup! I concur